# Contrastive Alignment with Semantic Gap-Aware Corrections in Text-Video Retrieval

### v4

@@ B.3 Ablation on $\Delta$ Norm Regularization Strength
$\mathcal{L}_{\varepsilon-\text{LogSumExp}}=\log\left(1+\frac{1}{B}\sum_{j=1}^{B}\exp\left(-\left(\varepsilon_{ij}-\overline{\varepsilon}_{i}\right)^{2}\right)\right),\quad\overline{\varepsilon}_{i}=\frac{1}{B}\sum_{j=1}^{B}\varepsilon_{ij}$
@@
Although this variant imposes a natural lower bound and provides vanishing gradients near convergence, it does not significantly improve retrieval. We hypothesize that this is due to gradient saturation when the variance among $\varepsilon_{ij}$ becomes too large, which in turn weakens the ability to further enforce directional separation. Unlike contrastive margins that encourage explicit pairwise gaps, the log-sum-exp loss lacks sufficient directional tension to align with the InfoNCE objective.
B.4 Effect of Anchor Choice in $\mathcal{L}_{\text{IB}}$
@@
Table 6: Effect of anchor choice in KL regularization $\mathcal{L}_{\text{IB}}$. Anchoring on $v$ yields better performance.
Anchor R@1 R@5 R@10 MedR MnR
$t_{i}$ (text) 47.9 74.9 83.8 2.0 12.5
$v_{j}$ (video) 49.1 74.7 83.6 2.0 12.0
To examine the role of modality anchoring in the information bottleneck regularization, we compare two variants of the KL divergence loss $\mathcal{L}_{\text{IB}}$ applied on the learned increments $\Delta_{ij}$. In both cases, $\Delta_{ij}$ is modeled as a Gaussian variable, but we anchor its distribution either from the video side (i.e., conditionally modeling $\Delta$ given $v_{j}$) or the text side ($t_{i}$). Results are shown in Table 6.
We observe that using the video as the anchor yields significantly better retrieval performance. This suggests that the information bottleneck not only suppresses unnecessary variance in $\Delta$, but also guides the model to extract task-relevant signals from the noisier modality—videos in this case. Since $\Delta$ is generated based on the initial pairwise difference ($v_{j}-t_{i}$ or $t_{i}-v_{j}$), it inherently reflects the model’s prior on how to reconcile the cross-modal gap. Videos tend to carry more diverse and redundant visual information, and anchoring the KL constraint on $v_{j}$ allows the model to compress this input through $\Delta$, effectively learning a semantically meaningful adjustment. Conversely, anchoring on $t_{i}$ fails to suppress visual redundancy, and the resulting $\Delta$ captures less structured or discriminative transformations, thereby degrading alignment quality. This analysis highlights that $\Delta$ acts not only as a tension offloader and false-negative noise suppressor, but also as a structured bottleneck pathway between text and video.
Appendix C Analysis of $\Delta$ Behavior During Training
@@
To better understand the role and dynamics of the learned modality-gap vector $\Delta$ throughout training, we visualize four key aspects of $\Delta$’s behavior, presented in Figure A1 to A4. These analyses reveal the underlying geometric transformations and provide further insight into how $\Delta$ facilitates optimization under InfoNCE loss. We will continue conducting ablation studies on the Text-to-Video retrieval task using the MSR-VTT [45] dataset.
C.1 Angle between $\Delta$ and $(v_{j}-t_{i})$
@@
(d) Pair distances: $t_{i}$/$t_{\Delta}$ vs. $v_{j}$
@@
As shown in Figure 4(a), we track the angle between $\Delta$ and the initial cross-modal gap vector $v_{j}-t_{i}$ throughout training, for both positive and negative pairs. At initialization, this angle is close to $\frac{\pi}{2}$, indicating that $\Delta$ is nearly orthogonal to $v_{j}-t_{i}$, and thus carries no meaningful alignment with the cross-modal semantic gap. This implies that the early $\Delta$ vectors do not differentiate between positive and negative pairs, acting more like isotropic perturbations in space. As training proceeds, however, this angle gradually increases into the obtuse region for both positive and negative samples. This reflects a significant shift in behavior—rather than attempting to directly bridge the modality gap $v_{j}-t_{i}$, the model learns to push $t_{i}$ away from $v_{j}$, effectively offloading the gradient tension induced by the modality gap and false negative interference. This offloading allows contrastive learning to take place in an expanded embedding space, where $\Delta$ modulates the representation geometry to ease the optimization burden. This trend is corroborated by Figure 4(d), where the Euclidean distances between the updated text embeddings $t_{\Delta_{ij}}=t_{i}+\Delta_{ij}$ and $v_{j}$ become significantly larger than the original distances $\|t_{i}-v_{j}\|$, for both positive and negative samples. That is, $\Delta$ introduces a global scaling effect in the representation space, and contrastive optimization is carried out under a larger geometric regime.
@@
Moreover, in early training stages, the angles between $\Delta$ and $v_{j}-t_{i}$ are nearly identical for positive and negative samples, showing that the model has not yet learned to encode fine-grained pairwise differences. But as training advances, a slight but consistent gap emerges—$\Delta$ for positive samples tends to have slightly smaller angles than that of negatives. This subtle divergence signals that $\Delta$ has begun to capture semantically meaningful pair-level distinctions, enabling more discriminative alignment in later stages.
C.2 Angle between $\Delta$ and $t_{i}$
@@
However, over time, this alignment loosens—$\Delta$ deviates further from $t_{i}$ as the model begins to adapt to modality-specific differences. This marks a transition where the network $\psi$ no longer treats the anchor $t_{i}$ as a default direction and instead generates $\Delta$ based on nuanced distinctions across the visual features $v_{j}$, thus leading to more expressive and semantically grounded $\Delta$ vectors.
C.3 Pair Distances between Text and Video Features
@@
Appendix D Why We Don’t Use an Initial $\Delta_{ij}^{*}=0$ Expansion and the Need for a Non-Zero Initial $\Delta_{ij}^{*}$
D.1 Introduction to the Issue
In the previous approach, we considered using multiple variables $\Delta_{ij}$ to distribute the gradient impact of each video $v_{j}$ on the text embedding $t_{i}$. However, the specific form of $\Delta_{ij}$ was unclear. The idea was to update each $\Delta_{ij}$ in the direction that would reduce the InfoNCE loss. A natural first step was to expand the loss function $\mathcal{L}_{i}$ as a Taylor expansion around the initial state $\Delta_{ij}^{*}=0$, which would correspond to an expansion solely based on $t_{i}$.
However, this expansion introduces a key issue: the gradient terms in each $\Delta_{ij}$ expansion are computed assuming that all other $\Delta_{ik}$ are zero. This neglects the interactions between different $\Delta_{ij}$ and effectively decouples the updates for each pair. Since InfoNCE loss requires comparing relative similarities between different pairs, this decoupled approach violates the fundamental nature of the contrastive learning objective, as it does not respect the relative comparisons between the cosine similarities of different text-video pairs.
D.2 Why Expanding with $\Delta_{ij}^{*}=0$ Is Inadequate

### v5

@@ C.2 Ablation of the Scale Coefficient $\alpha$ and the Orthogonality Variant in Directional Diversity
This rigid treatment prevents $\Delta_{ij}$ from encoding opposing semantic relations, indicating that directional diversity should remain semantically flexible rather than uniformly orthogonal.
Nevertheless, the orthogonalization approach exhibits stronger performance on video-to-text retrieval (R@1=49.4, R@5=75.4, R@10=83.6, Mean R=8.2). We conjecture that this improvement stems from the fact that the directional diversity regularization is applied with text as the anchor over candidate videos, where enforcing orthogonality among $\Delta_{ij}$} encourages the model to better separate diverse visual counterparts for the same textual query.
C.3 Ablation on $\Delta$ Norm Regularization Strength
@@
$\mathcal{L}_{\varepsilon-\text{LSE}}=\log\Big[1+\frac{1}{B}\sum_{j=1}^{B}\exp\big[-(\varepsilon_{ij}-\overline{\varepsilon}_{i})^{2}\big]\Big].$
@@
Although this variant imposes a natural lower bound and provides vanishing gradients near convergence, it does not significantly improve retrieval. We hypothesize that this is due to gradient saturation when the variance among $\varepsilon_{ij}$ becomes too large, which in turn weakens the ability to further enforce directional separation.
C.4 Ablation on the Anchor Choice in Relaxed VIB Regularization
@@
Table 9: Effect of anchor choice in KL regularization $\mathcal{L}_{\text{IB}}^{\text{relax}}$. Anchoring on $v$ yields better performance.
Anchor R@1 R@5 R@10 MdR MnR
$t$ 47.9 74.9 83.8 2.0 12.5
$v$ 49.1 74.7 83.6 2.0 12.0
We further investigate the effect of the anchor choice in the relaxed information bottleneck term $\mathcal{L}_{\text{IB}}^{\text{relax}}$. In the main formulation, the KL regularization is anchored on videos, i.e.,
$\mathbb{E}_{(t,v)}[\mathrm{KL}(q_{\psi}(\Delta|t,v)\|r(\Delta))]\geq\mathbb{E}_{v}\big[\mathrm{KL}(\bar{q}_{\psi}(\Delta|v)\|r(\Delta))\big],$
where $\bar{q}_{\psi}(\Delta|v)=\mathbb{E}_{t|v}[q_{\psi}(\Delta|t,v)]$. This relaxation aggregates increments ${\Delta_{ij}}_{i=1}^{B}$ under each video anchor $v_{j}$ and preserves the information-bottleneck effect while reducing sensitivity to the variability of short text inputs.
To contrast this design, we also test the reversed relaxation that anchors on text:
$\mathbb{E}_{(t,v)}\![\mathrm{KL}\!(q_{\psi}(\Delta\mid t,v)\,\|\,r(\Delta))]\geq\mathbb{E}_{t}\!\big[\mathrm{KL}\!(\bar{q}_{\psi}(\Delta\mid t)\,\|\,r(\Delta))\big],$
where $\bar{q}_{\psi}(\Delta|t):=\mathbb{E}_{v|t}[q_{\psi}(\Delta|t,v)]$. Practically, this means that the set $\{\Delta_{ij}\}_{j=1}^{B}$ associated with each text anchor $t_{i}$ is modeled as a Gaussian posterior, yielding the corresponding compression loss. Notably, our relaxation introduces stochasticity directly from the dataset pairing itself, rather than by sampling additional uncertainty variables.
As shown in Table 9, anchoring the KL term on the video side leads to superior retrieval performance (R@1=49.1 vs. 47.9). We attribute this to the inherent data characteristics: the textual descriptions in the dataset are typically short and less informative, while videos are semantically richer and often correspond to multiple captions. Consequently, conditioning $\Delta$ on $v$ captures more shared, modality-invariant structure, aligning better with the underlying multimodal distribution and yielding stronger retrieval results.
Figure 6: Comparison of hard negative alignment before and after applying $\Delta_{ij}$ optimization. Compared with the baseline, GARE produces smaller similarity gaps among semantically related videos $v_{j}$. This indicates that GARE effectively mitigates the noise from hard negatives and reduces the semantic deviation of the anchor $t_{i}$, leading to more stable and consistent alignment across similar samples.
C.5 Hard Negative Comparison
We compare GARE with two recent methods that explicitly or implicitly handle hard negatives:
DMAE jiang2023dual (R@1: 46.9, +1.6% over base. ACM MM 2023): DMAE enhances fine-grained alignment by mining hard positives—for instance, text queries associated with specific frames—which implicitly improves the model’s capability to separate hard negatives. Conceptually, this shares similarity with our variational information bottleneck (IB) loss, where $\Delta$ is compressed through a bottleneck to retain critical alignment signals while filtering noisy gradients.
NeighborRetr lin2025neighborretr (R@1: 49.5, +2.3% over base. CVPR 2025): This method employs a memory bank to compute $k$-neighbor co-occurrence statistics, identifying “good hubs” that encourage local consistency and reduce over-penalization of hard negatives. Although not explicitly formulated as hard-negative mining, the top-$k$ co-occurrence selection serves a similar role by adaptively reweighting difficult negatives.
Figure 7: Mean and variance of total gradients acting on $t_{i}$ on each dimension.
GARE (R@1: 49.1, +2.5% over base): Unlike the above methods, GARE does not rely on explicit mining or external memory structures. Each increment $\Delta_{ij}$ absorbs loss gradients locally from its paired $(t_{i},v_{j})$ sample, mitigating the reliance on global contrastive comparisons when encountering hard negatives. This local reallocation of gradients softens noisy updates and improves generalization. As illustrated in Figure 6, the text anchor $t_{i}$ in GARE exhibits enhanced semantic stability: for semantically related videos $v_{j}$, the similarity values $s_{ii}$ and $s_{ij}$ vary smoothly, indicating that $\Delta_{ij}$ enables more accurate fine-grained alignment across similar samples. This aligns with the gradient visualization in Figure 7, where the gradients on $t_{i}$’s dimensions remain centered around zero after introducing $\Delta$, demonstrating its role as a stable semantic prototype.
Efficiency comparison.
GARE maintains superior efficiency despite comparable or higher performance. It uses only a single cross-attention layer, whereas NeighborRetr includes 8 MLP modules and multiple transformer or convolutional blocks. NeighborRetr also requires large-scale memory banks (10,240 samples per modality) and $\sim$4.5 h training time, while GARE achieves similar accuracy with 1 h 34 min of training and negligible additional memory cost compared to the baseline. Together, these results demonstrate that GARE achieves comparable hard-negative robustness with substantially lower computational overhead.
Appendix D Derivation of the VIB Objective for Pair-Specific Increments
We model each pair-specific increment as a latent variable $Z=\Delta_{ij}$ for the input $X=(t_{i},v_{j})$ and the match label $Y\in\{0,1\}$. Our objective maximizes predictive information under a compression constraint:
$\max_{p(z|x)}I(Z;Y)-\beta\,I(Z;X),\quad\beta>0.$ (16)
Below we derive a tractable variational lower bound of Eq.  (16).
Lower bound for $I(\Delta_{ij};Y)$.
For convenience, in the following, we will refer to $y_{ij}$, $\Delta_{ij}$, $t_{i}$, and $v_{j}$ as $y$, $\Delta$, $t$, and $v$ respectively. We make the standard assumption that the match label does not depend on $\Delta$ once $(t,v)$ is given:
$p(y|\Delta,t,v)\;=\;p(y|t,v).$ (17)
By definition,
$I(\Delta;y)$ $=\iint p(\Delta,y)\log\frac{p(y|\Delta)}{p(y)}\,d\Delta dy$ (18)
$=\iint p(\Delta,y)\log\!\Big(\frac{q_{\theta}(y|\Delta)}{p(y)}\cdot\frac{p(y|\Delta)}{q_{\theta}(y|\Delta)}\Big)\,d\Delta dy$
$=\iint p(\Delta,y)\log\frac{q_{\theta}(y|\Delta)}{p(y)}\,d\Delta dy+\mathrm{KL}\big(p(y|\Delta)\|q_{\theta}(y|\Delta)\big)$
$\geq\iint p(\Delta,y)\log\frac{q_{\theta}(y|\Delta)}{p(y)}\,d\Delta dy\quad\text{(by non-negativity of KL)}$
$=\iint p(\Delta,y)\log q_{\theta}(y|\Delta)\,d\Delta dy+H(Y)$
$\geq\iiint p(y,\Delta,t,v)\log q_{\theta}(y|\Delta)\,d\Delta dyd(t,v)$
$=\iiint p(y|t,v)p(\Delta|t,v)p(t,v)\log q_{\theta}(y|\Delta)\,d\Delta dyd(t,v)$
$=\mathbb{E}_{(t,v,y)}\mathbb{E}_{\Delta\sim q_{\psi}(\Delta|t,v)}\!\big[\log q_{\theta}(y|\Delta)\big].$
where $q_{\theta}(y|\Delta)$ is a variational classifier, the non-negativivty entropy term $H(Y)$ is constant w.r.t. model parameters, and the last second equality uses assumption Eq. (17). Let $q_{\psi}(\Delta|t,v)$ be the variational encoder. Then,
$I(\Delta;y)\geq\mathbb{E}_{(t,v,y)}\,\mathbb{E}_{\Delta\sim q_{\psi}(\Delta\mid t,v)}\!\big[\log q_{\theta}(y|\Delta)\big].$ (19)
The two inequalities above arise from the non-negativity of $\mathrm{KL}(\cdot\|\cdot)$ and from rewriting $\log\frac{1}{p(y)}$ as the entropy $H(y)$.
Upper bound for $I(\Delta_{ij};\,t_{i},v_{j})$.
Using $I(\Delta;\,t,v)\leq\mathbb{E}_{(t,v)}\mathrm{KL}(p(\Delta|t,v)\|r(\Delta))$ with any auxiliary prior $r(z)$,
$I(\Delta;t,v)$ $=\iint p(\Delta,t,v)\log\frac{p(\Delta,t,v)}{p(\Delta)\,p(t,v)}\,d\Delta d(t,v)$ (20)
$=\iint p(\Delta,t,v)\log\frac{p(\Delta|t,v)}{p(\Delta)}\,d\Delta d(t,v)$
$=\iint p(\Delta,t,v)\log\Big(\frac{p(\Delta|t,v)}{r(\Delta)}\cdot\frac{r(\Delta)}{p(\Delta)}\Big)\,d\Delta d(t,v)$
$=\iint p(\Delta,t,v)\log\frac{p(\Delta|t,v)}{r(\Delta)}\,d\Delta d(t,v)-\mathbb{E}_{(t,v)}\mathrm{KL}(p(\Delta)\|r(\Delta))$
$\leq\iint p(\Delta,t,v)\log\frac{p(\Delta|t,v)}{r(\Delta)}\,d\Delta d(t,v)$
$=\mathbb{E}_{(t,v)}\big[\mathrm{KL}(p(\Delta|t,v)\|r(\Delta))\big]$
where the inequality follows from $\mathrm{KL}(p(\Delta)\|r(\Delta))\geq 0$ and $p(\Delta|t,v)$ denotes the variational encoder $q_{\psi}(\Delta|t,v)$ that parameterizes the model’s conditional distribution of the pair-specific increment (deterministic in our implementation). We use a diagonal Gaussian prior $r(z)=\mathcal{N}(0,I)$.
VIB lower bound.
Combining Eq. (18) and Eq. (20) yields the variational lower bound of Eq. (16):
$\mathcal{L}_{\text{VIB}}:=-\mathbb{E}_{(t,v,y)}\mathbb{E}_{\Delta\sim q_{\psi}(\Delta|t,v)}\big[\log q_{\theta}(y|\Delta)\big]+\beta\cdot\mathbb{E}_{(t,v)}\big[\mathrm{KL}(q_{\psi}(\Delta|t,v)\|r(\Delta))\big].$ (21)
Appendix E Relation between the Relaxed VIB Compression Term and the Original IB Objective
We start from the standard information bottleneck (IB) formulation, which regularizes the mutual information between the learned increment $\Delta$ and the input pair $(t,v)$:
$I(\Delta;T,V)=\mathbb{E}_{p(t,v)}\!\left[\mathrm{KL}\!\left(q_{\psi}(\Delta\mid t,v)\,\|\,q_{\psi}(\Delta)\right)\right].$ (22)
Since the marginal $q_{\psi}(\Delta)$ is intractable, it is commonly replaced by a fixed prior $r(\Delta)$, leading to the following upper bound:
$\mathbb{E}_{(t,v)}\!\left[\mathrm{KL}\!\left(q_{\psi}(\Delta\mid t,v)\,\|\,r(\Delta)\right)\right]=I(\Delta;T,V)+\mathrm{KL}\!\left(q_{\psi}(\Delta)\,\|\,r(\Delta)\right).$ (23)
This term penalizes the total amount of information $\Delta$ retains about both modalities and serves as the compression loss in the original VIB objective.
Relaxation anchored on the video side. To mitigate sensitivity to text-side variability, we anchor the expectation on videos and aggregate increments over all textual pairs associated with the same $v$. By convexity of the KL divergence (Jensen’s inequality), we have:
$\mathbb{E}_{(t,v)}\!\left[\mathrm{KL}\!\left(q_{\psi}(\Delta\mid t,v)\,\|\,r(\Delta)\right)\right]$ $=\mathbb{E}_{v}\mathbb{E}_{t\mid v}\!\left[\mathrm{KL}\!\left(q_{\psi}(\Delta\mid t,v)\,\|\,r(\Delta)\right)\right]$ (24)
$\geq\mathbb{E}_{v}\!\left[\mathrm{KL}\!\left(\bar{q}_{\psi}(\Delta\mid v)\,\|\,r(\Delta)\right)\right],$
where the aggregated posterior $\bar{q}_{\psi}(\Delta\mid v)=\mathbb{E}_{t\mid v}[q_{\psi}(\Delta\mid t,v)]$ represents the mixture distribution of increments under the same video anchor.
Applying the decomposition in Eq. (23) to the right-hand side of Eq. (24) yields
$\mathbb{E}_{v}\!\left[\mathrm{KL}\!\left(\bar{q}_{\psi}(\Delta\mid v)\,\|\,r(\Delta)\right)\right]=I_{\bar{q}}(\Delta;V)+\mathrm{KL}\!\left(q_{\psi}(\Delta)\,\|\,r(\Delta)\right),$ (25)
where $I_{\bar{q}}(\Delta;V)$ denotes the mutual information defined under the aggregated distribution $\bar{q}_{\psi}$. Combining the two equations gives
$I(\Delta;T,V)\;\geq\;I_{\bar{q}}(\Delta;V),$ (26)
indicating that the relaxation effectively removes the conditional term $I(\Delta;T\mid V)$ from the full IB regularization.
Interpretation. The relaxed KL term therefore optimizes
$\mathcal{L}_{\text{IB}}^{\text{relax}}=\mathbb{E}_{v}\!\left[\mathrm{KL}\!\left(\bar{q}_{\psi}(\Delta\mid v)\,\|\,r(\Delta)\right)\right]=I_{\bar{q}}(\Delta;V)+\mathrm{KL}\!\left(q_{\psi}(\Delta)\,\|\,r(\Delta)\right),$ (27)
which serves as a lower bound of the original compression term. While the standard IB loss penalizes all information in $\Delta$ about $(T,V)$, the relaxed form only constrains information shared with $V$ and ignores the conditional mutual information $I(\Delta;T\mid V)$. Intuitively, this relaxation preserves the bottleneck effect but allows $\Delta$ to encode text-specific variations within each video cluster, yielding more stable optimization when multiple captions correspond to the same visual content.
Appendix F Overall Objective and Inference
The overall training objective is formulated as
$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{info}}+\beta\,\mathcal{L}_{\text{IB}}^{\text{relax}}}_{\mathcal{L}_{\text{VIB}}}+\lambda_{\varepsilon}\,\mathcal{L}_{\varepsilon}+\lambda_{\text{dir}}\,\mathcal{L}_{\text{dir}},$ (28)
where the first two terms constitute the VIB optimization objective, and $\lambda{\varepsilon}$ and $\lambda_{\text{dir}}$ are the weights of the structural regularizers. In practice, we set $\beta=0.07$ and $\lambda_{\varepsilon}=\lambda_{\text{dir}}=0.01$ for MSR-VTT. During inference, we retain the learned increment $\Delta$ to assist retrieval, as it encodes the semantic consistency between $t_{i}$ and $v_{j}$. We further discuss the efficiency of using $\Delta$ at both training and inference stages in the Appendix I, as well as its distinction from traditional dual-encoder retrieval paradigms and deployment strategies for large-scale retrieval.
Appendix G Analysis of $\Delta$ Behavior During Training
@@
To better understand the role and dynamics of the learned semantic-gap vector $\Delta$ throughout training, we visualize four key aspects of $\Delta$’s behavior, presented in Figure 8(a) to 8(d). These analyses reveal the underlying geometric transformations and provide further insight into how $\Delta$ facilitates optimization under InfoNCE loss. We will continue conducting ablation studies on the Text-to-Video retrieval task using the MSR-VTT xu2016msr dataset.
@@
(d) Pair distances: $t_{i}$$t_{\Delta}$ vs. $v_{j}$
@@
Angle between $\Delta$ and $(v_{j}-t_{i})$.
As shown in Figure 8(a), we track the angle between $\Delta$ and the initial cross-modal gap vector $v_{j}-t_{i}$ throughout training, for both positive and negative pairs. At initialization, this angle is close to $\frac{\pi}{2}$, indicating that $\Delta$ is nearly orthogonal to $v_{j}-t_{i}$, and thus carries no meaningful alignment with the cross-modal semantic gap. This implies that the early $\Delta$ vectors do not differentiate between positive and negative pairs, acting more like isotropic perturbations in space. As training proceeds, however, this angle gradually increases into the obtuse region for both positive and negative samples. This reflects a significant shift in behavior—rather than attempting to directly bridge the semantic gap $v_{j}-t_{i}$, the model learns to push $t_{i}$ away from $v_{j}$, effectively offloading the gradient tension induced by the modality gap and false negative interference. This offloading allows contrastive learning to take place in an expanded embedding space, where $\Delta$ modulates the representation geometry to ease the optimization burden. This trend is corroborated by Figure 8(d), where the Euclidean distances between the updated text embeddings $t_{\Delta_{ij}}=t_{i}+\Delta_{ij}$ and $v_{j}$ become significantly larger than the original distances $\|t_{i}-v_{j}\|$, for both positive and negative samples. That is, $\Delta$ introduces a global scaling effect in the representation space, and contrastive optimization is carried out under a larger geometric regime.
@@
[…]
Angle between $\Delta$ and $t_{i}$.
@@
[…]
Pair distances between text and video features.
@@
Appendix H Why We Don’t Use an Initial $\Delta_{i*}=0$ Expansion and the Need for a Non-Zero Initial $\Delta_{i*}$
H.1 Introduction to the Issue
In the previous approach, we considered using multiple variables $\Delta_{ij}$ to distribute the gradient impact of each video $v_{j}$ on the text embedding $t_{i}$. However, the specific form of $\Delta_{ij}$ was unclear. The idea was to update each $\Delta_{ij}$ in the direction that would reduce the InfoNCE loss. A natural first step was to expand the loss function $\mathcal{L}_{i}$ as a Taylor expansion around the initial state $\Delta_{i*}=0$, which corresponds to an expansion solely based on $t_{i}$.
However, this expansion introduces a key issue: the gradient terms in each $\Delta_{ij}$ expansion are computed under the assumption that all other $\Delta_{ik}$ remain zero. This neglects the mutual interactions among $\{\Delta_{ij}\}_{j=1}^{B}$ and decouples the updates for different pairs. Since the InfoNCE loss fundamentally depends on the relative ranking of cosine similarities between $t_{i}$ and all candidate videos $\{v_{j}\}$, such a decoupled formulation disrupts the prior ordering structure already established by the CLIP encoder’s logits. In other words, it breaks the model’s inherent relational prior among candidates, which is essential for maintaining consistent contrastive comparisons.
H.2 Why Expanding with $\Delta_{i*}=0$ Is Inadequate